OpenStudy (anonymous):

solving equation arccos (2x^2-1) =2arccos (1/2)

OpenStudy (mertsj):

On the right...arccos(1?2)=pi/3. Do you agree?

OpenStudy (mertsj):

So now we have: arccos(2x^2-1)=2pi/3

OpenStudy (mertsj):

And not take cosine of both sides.

OpenStudy (mertsj):

2x^2-1=-1/2 Solve that equation

OpenStudy (mertsj):

And not take cosine of both sides. now....instead of not

OpenStudy (anonymous):

ok thanks Can I solve it in another way?

OpenStudy (mertsj):

Maybe but I do not know another method. Perhaps you could show me.

OpenStudy (anonymous):

I can not, I'm just wondering